A proton is accelerated from rest through a potent
Answer #1

Here

Potential , V = 1 kV = 1000 V

magnetic fiedld , B = 4.5 mT

B = 0.0045 T

a) for the redius

radius of path = sqrt(2 * m * e * V)/(B * e)

radius of path = sqrt(2 * 1.67 *10^-27 * 1.602 *10^-19 * 1000)/(0.0045 * 1.602 *10^-19 )

radius of path = 1.015 m

the radius of path is 1.015 m

b)

for the time period

time period = 2*pi * m/(B * q)

time period = 2pi *1.67 *10^-27/(0.0045 * 1.602 *10^-19)

time period = 1.46 *10^-5 s

the time period of each revolution is 1.46 *10^-5 s